OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions. Through (6, 7); parallel to the line passing through (7, 5) and (3, 1)

OpenStudy (anonymous):

First find the slope of the line since you know it is parallel to the line (7,5) and (3,1) you can simply find the slope for that line which is m=(Y2-Y1)/X2-X1)=4/4=1 now use point slope formula Y-Y1=m(X-X1) Y-7=X-6

OpenStudy (anonymous):

slope of the given parallel = 1 ====> y=mx+b ===> y=x+b ==> 7=6+b ==> b=1 ==> the required equation is y = x+1

OpenStudy (anonymous):

ooh okay...i just didnt know the first step..i think i can figure it out from there, thank you! if you have time i had a previous question about average velocity that i am stumped on haha...thanks guys